OpenStudy (anonymous):

hey there i have difficulty with augmented matrix ... can anybody help me with that?

OpenStudy (anonymous):

sure, what's up?

OpenStudy (anonymous):

system: x-4y+3z=5 -x+3y- z=-3 2x -4z=6

OpenStudy (anonymous):

1 3 2 / 1 0 1 -2 /5 0 0 4 /4

OpenStudy (anonymous):

not is not a system

OpenStudy (anonymous):

coefficient matrix 1 -4 3 -1 3 -1 2 0 -4

OpenStudy (anonymous):

augmented matrix 1 -4 3 5 -1 3 -1 -3 2 0 -4 6

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

no, that's an example, I'm not sure what your question is. you just typed in a matrix.

OpenStudy (anonymous):

your augmented matrix is derived from the coefficients and constant terms of your system of linear equations

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

do youwant me to write u the whole question maybe u gonna understant ??

OpenStudy (anonymous):

yes. are you given a system of linear equations & asked to write an augmented matrix?

OpenStudy (anonymous):

find X, where x=(x1,x2,x3)to the power of t, given the A is known (3*) matrix, b a known column matrix, and that the system Ax=b has augmented matrix equivalent to

OpenStudy (anonymous):

1 3 2 ! -1 0 1 -2 ! 5 0 0 4 ! 4

OpenStudy (anonymous):

thats is the qustion

OpenStudy (anonymous):

question sorry

OpenStudy (anonymous):

haha, yeah, sorry mattheos, that is a little more in depth than just having difficulty with an augmented matrix. i'm in my first semester of linear algebra & i haven't seen a problem like that yet.

OpenStudy (anonymous):

aham ok ... thx anyway

OpenStudy (anonymous):

I'm going to repost your question if you don't mind, I would like to see a solution as well, I could solve this if it didn't involve the power of t.

OpenStudy (anonymous):

Do you have to use gaussian elimination?